Okay so I'm trying to figure out how to evaluate this expression. 9^-3 • 9^5 over 9^-2 all to the negative half power.
Answer by Boreal(15235)

9^-3*9^5=9^2.
9^2/9^(-2)=9^4
(9^4)^(-1/2)=9^(-2)
=(1/81)